The first term of a geometric progression is √ 3+1 and the sum of the first three

terms is 3( √ 3-1) . Find the common ratio of the progression.

(4marks).


1
Expert's answer
2021-10-05T10:58:59-0400
"a=\sqrt{3}-1"

"a+ar+ar^2=3(\sqrt{3}-1)"

"a(1+r+r^2)=3(\sqrt{3}-1)"

"(\sqrt{3}-1)(1+r+r^2)=3(\sqrt{3}-1)"

"1+r+r^2=3"

"r^2+r-2=0"

"r(r-1)+2(r-1)=0"

"(r+2)(r-1)=0"

"r_1=-2, r_2=1"

The common ratio of the progression is "-2" or the common ratio of the progression is "1."


If we take that "r\not=0, r\not=1," then the common ratio of the progression is "-2".